John in conducting a two-sample t-test. In his data, Group 1 has 32 participants and Group 2 has 30 participants. What are the degrees of freedom for his test?
Answer #1

Solution::-

According to the given data John conducting a two sample t-test in which two groups are present

Group 1 has 32 participants

Group 2 has 30 participants

since they are conducting two-sample t-test ,

Relation to find degrees of freedom (df )= n1 + n2 -2

df=32+30-2

=62-2

=60

therefore the degrees of freedom (df)=60.